londonMarie:

f a sprinkler waters 1 over 15 of a lawn in 1 over 5 of an hour, how much time will it take to water the entire lawn? 2 hours 3 hours 10 hours 15 hours

Hoodmemes:

Ight so a whole, will be 15/15 or 1 whole,and the sprinkler does 1/15 of the lawn in 1/3 of an hour, how long will it be to do the whole 15/15?

Hoodmemes:

lawn hours 1/15 15/15 1/3 h

Hoodmemes:

so \[\frac{ 1 }{ 15 } \frac{ 15 }{ 15}=\frac{ 1 }{ 3 } h \rightarrow \frac{ 1 }{ 15 }\times \frac{ 15 }{ 15}\]

Hoodmemes:

1\[\frac{ 1 }{ 15 }\times\frac{ 15 }{ 15 }=\frac{ 1 }{ 3 }\frac{ 1 }{ }\]

Hoodmemes:

*\[\frac{ 1 }{ 3 }\frac{ 1 }{ h }\]

Hoodmemes:

so \[\rightarrow \frac{ 1 }{ 15}=\frac{ 1 }{ 3h }=15\rightarrow 3h=15\rightarrow h=\frac{ 15 }{ 3 }\]\rightarrow h= (blank)

Hoodmemes:

\[\frac{ 15 }{ 3}\rightarrow h=(blank)\]

Hoodmemes:

ima hope dis helps

londonMarie:

thx

Hoodmemes:

np